Suppose that a die is rolled twice. What are the possible values and their corresponding probabilities that the following random variables can take on: (a) the maximum value to appear in both rolls; (b) the sum of the values obtained from the two rolls; (c) the value of the first roll minus the value of the second roll?


1
Expert's answer
2021-03-01T16:08:56-0500

When dies is rolled twice The sample space are-




(a) The maximum value appears in both rolls is "(6,6)" .

and probability "=" "\dfrac{1}{36}"


(b) The sum of the values obtained from rolls are-

"2\rightarrow (1,1)\\3\rightarrow(2,1),(1,2)\\4\rightarrow (1,3),(2,2),(3,1)\\5\rightarrow (1,4),(2,3),(3,2),(4,1)\\6\rightarrow (1,5),(2,4),(3,3),(4,2),(5,1)\\7\rightarrow (1,6),(2,5),(3,4),(4,3),(5,2),(6,1)\\8\rightarrow (2,6),(3,5),(4,4),(5,3),(6,2)\\9\rightarrow(3,6),(4,5),(5,4),(6,3)\\10\rightarrow(4,6),(5,5),(6,4)\\11\rightarrow(5,6),(6,5)\\12\rightarrow(6,6)"


Probability that sum is "2=\dfrac{1}{36}"


Probability that sum is "3=\dfrac{2}{36}"


Probability that sum is "4=\dfrac{3}{36}"


Probability that sum is "5=\dfrac{4}{36}"


Probability that sum is "6=\dfrac{5}{36}"


Probability that sum is "7=\dfrac{6}{36}"


Probability that sum is "8=\dfrac{5}{36}"


Probability that sum is "9=\dfrac{4}{36}"


Probability that sum is "10=\dfrac{3}{36}"


Probability that sum is "11=\dfrac{2}{36}"


Probability that sum is "12=\dfrac{1}{36}"


(c) Value of the first roll minus the value of the second roll can hev the values 0,1,2,3,4,5


Probabilty that diffirence is "0=\dfrac{6}{36}"


Probabilty that diffirence is "1=\dfrac{10}{36}"


Probabilty that diffirence is "2=\dfrac{8}{36}"

Probabilty that diffirence is "3=\dfrac{6}{36}"


Probabilty that diffirence is "4=\dfrac{4}{36}"


Probabilty that diffirence is "5=\dfrac{2}{36}"